# Scene: Procedural Controllable Large-scale Scene Generation via Large-language Models

### v1

@@ Abstract
Due to its great application potential, large-scale scene generation has drawn extensive attention in academia and industry. Recent research employs powerful generative models to create desired scenes and achieves promising results. However, most of these methods represent the scene using 3D primitives (e.g. point cloud or radiance field) incompatible with the industrial pipeline, which leads to a substantial gap between academic research and industrial deployment. Procedural Controllable Generation (PCG) is an efficient technique for creating scalable and high-quality assets, but it is unfriendly for ordinary users as it demands profound domain expertise. To address these issues, we resort to using the large language model (LLM) to drive the procedural modeling. In this paper, we introduce a large-scale scene generation framework, Scene, which can automatically produce high-quality procedural models according to designers’ textual descriptions.Specifically, the proposed method comprises two components, PCGBench and PCGPlanner. The former encompasses an extensive collection of accessible procedural assets and thousands of hand-craft API documents. The latter aims to generate executable actions for Blender to produce controllable and precise 3D assets guided by the user’s instructions. Our Scene can generate a city spanning 2.5 km 2.5 km with delicate layout and geometric structures, drastically reducing the time cost from several weeks for professional PCG engineers to just a few hours for an ordinary user. Extensive experiments demonstrated the capability of our method in controllable large-scale scene generation and editing, including asset placement and season translation. Our project page: https://scenex-lab.github.io/.
Keywords:
Large-scale Scene Generation LLM Agents PCG
1 Introduction
With the rapid progress in generative modeling, 3D content generation has witnessed remarkable achievements in automatic creation of 3D objects [39, 43, 35, 36], avatars [58, 27, 17], and scenes [31, 52, 29, 9]. In this work, we focus on automatic 3D scene generation, which has extensive applications in domains like gaming, animation, film, and virtual reality. Although previous approaches have achieved promising results, a significant gap still remains for industrial-grade applications due to incompatible representations with the industrial pipeline of the scene using 3D primitives (e.g. point cloud or radiance field).
To bridge the gap, we resort to procedural modeling, an advanced technique for the fast creation of high-quality 3D assets. It can generate realistic and intricate 3D assets in natural scenes through adjustable parameters and rule-based systems [32] using software like Blender. For example, Infinigen [42] proposes a procedural generator that combines different procedural modeling algorithms to generate large-scale natural scenes encompassing terrain, weather, vegetation, and wildlife. [33], [5] and [46] use procedural modeling to generate city-level street or layout. These procedural approaches generate high-quality 3D assets at the object, scene, and city levels. However, the comprehensive grasp of generation rules, algorithmic frameworks, and individual parameters of procedural modeling make the generation process beginner-unfriendly and time-consuming, especially for large-scale scene or city generation. For instance, generating a New York city, as shown in Fig 1, requires the effort of a professional PCG engineer working for over two weeks. Moreover, due to the substantial learning curve for users, it is formidable to generate large-scale scenes using procedural modeling.
To address the above problems, existing works such as 3D-GPT [45] and SceneCraft [19] introduce an instruction-driven 3D modeling method by integrating an LLM agent with procedural generation software. Despite the successful collaboration with human designers to establish a procedural generation framework, these methods still exhibit significant limitations. Firstly, SceneCraft is unable to achieve fine-grained 3D asset editing due to the uneditable nature of fixed 3D assets compared to those created by procedural generation. Secondly, 3D-GPT, being based on the parametric procedural generation model Infinigen, restricts its ability to fully utilize existing PCG resources. Additionally, it is incapable of generating high-quality and large-scale scenes, especially expansive terrains and boundless cities. Lastly, the diversity of assets is limited due to constraints imposed by prompt expressions.
To address these issues, we propose an instructive large-scale scene generation framework, Scene, which can automatically generate high-quality scenes according to designers’ textual descriptions. By introducing a PCGBench, we collect extensive accessible procedural assets, including various terrains, cities, and natural objects like plants and animals. This helps us to generate diverse, realistic, and practical large-scale scenes. Moreover, we also construct thousands of API parameter documents, which are used to formulate the agent planner process and provide a quantitative evaluation. Inspired by existing LLM-driven planning methods [49, 44, 51, 28], we develop a PCGPlanner framework to produce correct API actions interfaced with Blender, to automatically generate controllable and precise 3D assets guided by users’ instructions. This LLM-based automated solution significantly enhances the efficiency of scene generation, reduces the learning curve for procedural scene modeling, and enables the customized creation of target scenes. In addition to the inherent advantages of procedural modeling such as photorealistic and real geometry, our Scene possesses several key properties:
@@ 1 Introduction
Efficiency: Benefiting from the proposed PCGBench and PCGPlanner, our Scene can rapidly produces extensive, high-quality 3D assets, including terrain, city, and forest. Moreover, we only need a few hours to generate a large-scale city, whereas it would take a professional designer over two weeks.
@@
Controllability: Scene can generate 3D scenes satisfying personalized demands. For example, we can achieve scene editing according to the corresponding descriptions, such as adding objects, placing objects at a location, translating the season, and so on.
@@
Diversity: Scene breaks the limitations of specific generation algorithms and predefined rules, leveraging advanced language model technology to achieve more flexible and diverse large-scale scene generation. Specifically, we generate large-scale terrains in the natural world, like deserts, lakes, forests, and mountains, and high-quality cities in the real world, such as New York, Paris, and Tokyo, as well as various objects, including plants, animals, and buildings.
2 Related Works
2.1 Learning Based 3D Generation.
@@ 2 Related Works
In recent years, 3D asset generation has witnessed rapid progress, combining the ideas of computer graphics and computer vision to realize the free creation of 3D content. Presently, predominant research in 3D asset generation primarily concentrate on creating individual objects [39, 35, 30, 8], 3D avatars [58, 27, 18, 25], and 3D scenes [16, 61, 11, 62, 59]. Among these, ZeroShot123 [35] proposes a method based on a diffusion model, which realizes the 3D model of the target based on a picture. DreamFusion [39] proposes a NeRF-based approach that allows the model to generate a corresponding 3D model based on the input text.
Compared with the generation of a single object, the generation of large-scale scenes, including the generation of natural landforms [15, 34, 9] and borderless cities [31, 29, 52], is undoubtedly more practical and challenging in the metaverse and game production. CityDreamer [52] builds vast, large-scale 3D cities based on the layout of real cities, enhancing the accuracy and stability of urban reconstruction. SceneDreamer [9] proposes a method to generate 3D borderless scenes within 2D plots using BEV representations. These learning-based methods do not have geometric structures, making them difficult to apply directly to virtual engines such as UE. For example, although InfiniCity [31] and CityDreamer [52] can generate visually appealing urban scenes, the absence of clear geometry may lead to issues in practical applications, such as intersecting objects and spatial discontinuities.
2.2 Procedural Based 3D Generation.
@@
Researchers have delved into the procedural generation of natural scenes [12, 60] and urban scenes [33, 46, 48, 54] using Procedural Content Generation (PCG). For instance, PMC [38] proposes a procedural way to generate cities based on 2D ocean or city boundaries. It employs mathematical algorithms to generate blocks and streets and utilizes tailgating technology to generate the geometry of buildings. While these traditional computer graphic methods can produce high-quality 3D data, all parameters must be pre-entered into the procedurally generated process. The resultant 3D data is bound by rule limitations, introducing a certain level of deviation from the real world. This significantly constrains flexibility and practical usability. Infinigen [42] introduces a technique for procedurally generating realistic 3D natural objects and scenes. This methodology facilitates the programmatic generation of all assets, with meshes and textures generated through random mathematical algorithms. Although Infinigen generates infinitely real assets, users are unable to customize the generated outcomes because of their specific requirements. Therefore, we propose a more convenient procedural generation method to produce high-quality 3D assets.
2.3 LLM Based Task Planner
@@
Benefiting from knowledge hidden in the large-language model (LLMs) [41, 1, 10, 26, 3], researchers explore LLMs to address intricate tasks beyond canonical language processing domains. These tasks encompass areas such as mathematical reasoning [22, 50], medicine [53, 23], and planning [57, 14, 21, 20]. Thanks to powerful reasoning and generalization capabilities, LLMs act as practiced planners for different tasks. For example, [20] utilizes the expansive domain knowledge of LLMs on the internet and their emerging zero-shot planning capabilities to execute intricate task planning and reasoning. [14] investigates the application of LLMs in scenarios involving multi-agent coordination, covering a range of diverse task objectives. [56] presents a modular framework that employs structured dialogue through prompts among multiple large pretrained models. Moreover, specialized LLMs for particular applications have been explored, such as HuggingGPT [44] for vision perception tasks, VisualChatGPT [51] for multi-modality understanding, Voyager [49] and [63], SheetCopilot [28] for office software, and Codex [6] for Python code generation. Inspired by the existing works, we explore the LLM agent to PCG software, e.g., Blender, to provide automatic 3D assets generation.
3 Methods
The proposed Scene framework contains two main components: the PCGBench and PCGPlanner. PCGBench is a large-scale dataset that includes diverse resources in procedural modeling, which is used to perform as an asset bank and evaluate the generation performance. Specifically, it includes collecting plugins and assets from the Internet, asset annotations, and hand-craft API documents. Based on our PCGBench, a novel PCGPlanner framework is proposed to automatically generate 3D scenes using both PCG and pre-defined models. It comprises a task planner stage, an asset retrieval stage, and an action execution stage guided by agents with custom-built instructions. By the cooperation between the PCGBench and novel PCGPlanner agents, Scene is endowed with the capability of efficient, controllable generation of diverse, high-quality scenes.
3.1 PCGBench
To build PCGBench, we first collect abundant resources of PCG models from Blender Market111https://blendermarket.com/, which includes action APIs and 3D assets. The action APIs consist of raw APIs and detailed parameters that are used to control asset generation. Specifically, we have collected 1,532 individual raw APIs in total, involving each aspect in the scene-generation tasks, such as controlling generated assets’ placement, color, size, and height. Then, we manually merge the most essential raw APIs according to their functions and obtain 45 action functions. Each action function contains multiple raw APIs with the corresponding parameters to produce a complete 3D asset. For example, to generate a 3D tree as Fig 2 shows, we define two individual action functions such as tree_mesh_generate and tree_material_generate. Each function comprises several raw APIs with the corresponding parameters like TreeHight, CrownDensity, LeafDensity, etc. These parameters can be manipulated at the action input, and we can generate various trees by changing these parameters.
To achieve controllable 3D model generation guided by human instructions, we should formulate the action functions for large-language models (LLMs), e.g., GPT-4222https://openai.com/gpt-4, following previous works [49, 45, 6]. To achieve this, we compile lots of action functions guidance documents, denoted as API Documents, to deploy our functions for LLMs. Concretely, this document includes information on API types, function descriptions, API names, and a set of raw APIs with the corresponding parameter details. Besides, we also provide a comprehensive record of all parameter names and their possible values. The details are referred to Fig 2. Moreover, we have categorized these documents into four modules:{ Terrain, City, Detail, Weather}. Simultaneously, we have developed a concise document outlining the scene generation process, providing detailed explanations of the construction sequence and capability boundaries for each module.
Regarding asset collection, we have meticulously curated an extensive 3D dataset comprising 1,908 finely crafted models and 1,294 textures. This diverse collection spans categories including buildings, flora, urban infrastructure, and textures gathered from the vast expanse of the Internet. These existing assets provide an alternative solution for 3D scene generation by retrieving existing individual procedural models rather than progressively generation, which can significantly improve the efficiency of creating large-scale scenes. For retrieval, we use CLIP [40] encoder feature, achieving text-to-text retrieval and text-to-image retrieval. Therefore, we also provide abundant annotations for collected 3D assets by rendering a picture for each 3D asset.
3.2 PCGPlanner
To achieve LLM-driven asset generation, we introduce PCG Planner Agents’ hierarchy, which is comprised of multiple agents that can automatically generate high-quality and large-scale scenes corresponding to the users’ textual descriptions. Following previous works [44, 51, 49, 6], we design our module encompassing three primary stages: task planning, asset retrieval, and action execution. In the task planning stage, a sequential planning agent pair is introduced to process the user query and output a detailed sub-task list. Then, to precisely connect the sub-task with our PCGBench, we retrieve the actions and assets mentioned in the task list. Finally, the execution agent leverages the annotated action APIs to process subtasks of procedural generation or asset manipulation.
Agent construction To endow the LLMs with the capability for modeling scenes like a specialist, we introduce a systematic template to prompt LLMs as agents to achieve the subtasks in a scene modeling task. For each agent, the prompt has a similar structure defined as (, , , , ), where i {dispatch, specialist, retrieval, execution} distinguishes the responsibility of agents. The constituents of the prompt are defined as follows:
@@ 3.2 PCGPlanner
Document At each step, the agent is prompted by a knowledge document according to their task. We denote D as the collection that contains all the knowledge documentation pre-defined in PCGBench.
@@
Format F denoted as the output format for each agent. To precisely and concisely convey information between agents, the output format of each agent is strictly defined.
@@
Examples To help agents respond strictly following the format, we demonstrated several examples E for reference in each prompt.
3.2.1 Task Planner
For a scene generation task, multiple action APIs from different modules are usually needed. There is a great possibility that different APIs require to be called in a sequential order. Thus, a correct function arrangement between different modules and APIs is of essential importance. To ensure competence in handling a broad spectrum of scene-generation tasks and surmounting token limitations, we present a two-stage iterative task planner comprising a dispatch agent and a specialist agent as depicted in Fig. 3.
Dispatch agent Dispatch agent is the first agent to receive users’ descriptions of the target scene, it is responsible for extracting information from the sentences. It serves as a planner to interpret user intents into a rough plan and list all the objects mentioned in the input instruction. In the meantime, users may omit some common background objects in the description, the agent will automatically add all these objects that need to be taken into account in the rough plan.
The detailed object requirements of a scene and their module correspondence principle are described in the knowledge documentation . We formulate the dispatch agent as follows:
where is the user input, stands for an object listed in the rough plan in the formation of {Module: Object name}
Specialist agent Then, the specialist agent is activated to make further detailed plans.
At this stage, a specialist agent serves as a professional designer who is capable of giving detailed plans for the objects given in the rough plan. According to the module class of the object, the specialist agent is prompted with the corresponding knowledge documentation , where contains the generation mechanisms and explanations for each module. The specialist agent decomposes each rough plan into a detailed plan with several sub-tasks, where each subtask corresponds to an executable API or an asset to be imported
where stands for one of the sub-tasks to be completed.
3.2.2 Asset Retrieval
With the detailed plan given by the specialist agent, we need to search in our PCGBench for a suitable API to execute or an asset to import. To precisely retrieve them, we propose a retrieval agent that identifies whether the target of the sub-task type is an asset or an API. With the prompt , the agent is also instructed to output a description as the search index of the asset or the API. The description is then encoded into an embedding for the retrieval. Here, a pre-trained CLIP [40] is used to search our action APIs by text-to-text retrieval and 3D models by text-to-image retrieval, respectively. As shown in the right part of Fig 2, the rendering images of each asset in our PCGBench are encoded into normalized 768-dimensional vectors, and then compared with the input description embedding. One of the five most similar results is randomly selected as the outcome and imported into the Blender scene. The best-matched result is selected by calculating the similarity by cosine distance.
@@ 3.2.2 Asset Retrieval
where stands for an retrieved API, stands for the API collection. is the retrieved asset from asset collection . For text-to-text retrieval, the most relevant API is also selected by its functional description embedding. If the retrieval result is an asset, it will be directly imported into the scene project. On the other hand, the retrieved API is to be passed to the action execution agent.
3.2.3 Action Execution
In the PCGBench, each action API is defined in the form of a Python function with multiple pre-defined parameters. After obtaining the , we need to make custom modifications for these initial parameters to meet the task description. For each API , there is a knowledge documentation that passed into the prompt . As illustrated in Fig 2 (a), the knowledge documentation of each action API is recorded in a JSON format and has the same architecture. Take tree_mesh_generate API as an example, each parameter in the function is precisely defined and explained in tree_mesh_doc = {"type": "tree mesh", "Description": "Function for generating trees mesh.", "Function API": "tree_mesh_generate", "Parameter Information": {"Possible values": ("High", "Medium", "Low"), … } }. The execution agent can be expressed as:
Where is the manipulated version of retrieved API . There is a chance that the descriptions contain less information for all the parameters. In this situation, we instruct the execution agent to automatically choose a feasible parameter from the options according to the other parameters. The execution agent ultimately executed the action API with the inferred parameters in Blender Python environments, incrementally adding new objects and making modifications in scene project until all the tasks were completed.
4 Experiments
@@
The goals of our experiments are threefold: (i) to verify the capability of Scene for generating photorealistic large-scale scenes, including nature scenes and cities; (ii) to demonstrate the effectiveness of Scene for personalized editing, such as adding or changing; (iii) to compare different LLMs on the proposed benchmark.
4.1 Benchmark Protocol
@@
Metrics: We use Executability Rate (ER@1) and Success Rate (SR@1) to evaluate the capabilities of LLMs on our Scene. The former measures the proportion of proposed actions that can be executed, and the latter is used to evaluate action correctness [7]. Moreover, to quantify the aesthetic level of the generated scenes, we adopt a unified judgment standard as a reference. We divide the aesthetics of generated scenes into five standards: Poor (1-2 points)/Below Average (3-4 points)/Average (5-6 points)/Good (7-8 points)/Excellent (9-10 points). We enlisted 35 volunteers to assess the quality of our generation, including 5 PCG experts. We compute average score (AS) and average expert score (AES) to evaluate the effectiveness of our method.
4.2 Main Results
Large-Scale Scene Generation. To intuitively illustrate the capability of Scene in generating large-scale scenes, we select three subsets with significant differences for comparison. We perform quantitative and qualitative results, including generation quality, aesthetic evaluation and efficacy comparison.
Generation Quality. We first provide several examples of our Scene for generating large-scale nature scenes and unbounded cities. The results are shown in Fig 4. From the results, we can observe that the proposed Scene can produce highly realistic scenes in both natural scenes and cities. Moreover, the generated content is correctly corresponding to the provided texture descriptions. These demonstrate the power and effectiveness of our proposed LLM-driven automatic 3D scene generation framework.

### v3

@@ Abstract
Developing comprehensive explicit world models is crucial for understanding and simulating real-world scenarios. Recently, Procedural Controllable Generation (PCG) has gained significant attention in large-scale scene generation by enabling the creation of scalable, high-quality assets. However, PCG faces challenges such as limited modular diversity, high expertise requirements, and challenges in managing the diverse elements and structures in complex scenes. In this paper, we introduce a large-scale scene generation framework, Scene, which can automatically produce high-quality procedural models according to designers’ textual descriptions. Specifically, the proposed method comprises two components, PCGHub and PCGPlanner. The former encompasses an extensive collection of accessible procedural assets and thousands of hand-craft API documents to perform as a standard protocol for PCG controller. The latter aims to generate executable actions for Blender to produce controllable and precise 3D assets guided by the user’s instructions. Extensive experiments demonstrated the capability of our method in controllable large-scale scene generation, including nature scenes and unbounded cities, as well as scene editing such as asset placement and season translation. Our project page: https://zhouzq1.github.io/SceneX/.
1University of Chinese Academy of Sciences (UCAS), 2Centre for Artificial Intelligence and Robotics (CAIR)
3Institute of Automation, Chinese Academy of Sciences (CASIA), 4New Laboratory of Pattern Recognition (NLPR)
5Shandong University, 6State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS)
7University of Science and Technology Beijing, 8China University of Geosciences Beijing
{zhoumengqi2022, zhaoxiang.zhang}@ia.ac.cn,yuxiwang93@gmail.com,jrpeng4ever@126.com
Introduction
In the realm of Artificial General Intelligence (AGI), developing comprehensive world models is crucial for understanding and simulating real-world scenarios(Zhang et al. 2024d; Huang et al. 2024). Recent advancements, such as those demonstrated by the Sora model (Brooks et al. 2024), show progress in capturing physical laws and generating realistic simulations. However, Sora’s outputs often lack detailed geometry and structured information, limiting their editability and interactivity. To address these limitations, explicit world models provide a more robust solution. By constructing worlds with detailed mesh assets and evolving them according to predefined physical rules, these models leverage Physically Based Rendering (PBR) to ensure high geometric consistency and detailed visualization. Procedural modeling methods, such as those outlined by Lindenmayer et al. (Lindenmayer 1968), show great promise in generating realistic and intricate world models through adjustable parameters and rule-based systems using tools like Blender. For example, Infinigen (Raistrick et al. 2023) proposes a procedural generator to generate large-scale natural scenes encompassing terrain, weather, vegetation, and wildlife. (Lipp et al. 2011), (Chen et al. 2008) and (Talton et al. 2011) use procedural modeling to generate city-level street or layout. Although these procedural approaches generate high-quality 3D assets, they are beginner-unfriendly and time-consuming due to the comprehensive grasp of generation rules, algorithmic frameworks, and individual parameters of procedural modeling. For instance, generating a city, as shown in Fig. 1, requires the effort of a professional PCG engineer working for over two weeks.
To address the above problems, existing works such as 3D-GPT (Sun et al. 2023) and SceneCraft (Hu et al. 2024b) introduce an instruction-driven 3D modeling method by integrating an LLM agent with procedural generation software. Despite the successful collaboration with human designers to establish a procedural generation framework, these methods still exhibit significant limitations. Firstly, the restricted editing capabilities of fixed 3D resources prevent SceneCraft from attaining the same degree of precision in 3D resource editing as is achievable through procedural generation. Secondly, 3D-GPT is based on the procedural generation model Infinigen, which restricts its capacity to fully utilise existing PCG resources. This is exemplified by the inability to generate extensive terrain and limitless cities. Finally, the task of generating a large-scale scene is often composed of multiple related subtasks, such as the planning of the scene layout, the generation and placement of assets, and the adjustment of environmental details. However, there is currently no effective way to manage the execution order and dependencies of these tasks.
In this paper, we introduce the Scene framework, consisting of PCGHub and PCGPlanner. PCGHub addresses the limitations of individual PCG modules, which are constrained by their inherent algorithms and predefined rules. By integrating diverse procedural modules and encapsulating them with corresponding APIs, PCGHub offers a platform for continuously incorporating new PCG capabilities, thus supporting the expansion and diversity of procedural generation resources. Due to constraints in procedural algorithms and predefined rules, unrestricted combination of all procedural methods is infeasible. Thus, we develop the PCGPlanner to coordinate PCG methods within these constraints. PCGPlanner accesses and integrates procedural modules, enabling effective coordination and seamless integration. Our work improves scene generation flexibility and diversity, reduces entry barriers for non-coders, and efficiently uses existing technologies for accessible, community-friendly procedural generation. Our Scene possesses several key properties:
@@ Introduction
Efficiency: Benefiting from the proposed PCGHub and PCGPlanner, our Scene can rapidly produces extensive, high-quality 3D assets, including terrain, city, and forest. Moreover, we only need a few hours to generate a large-scale city, whereas it would take a professional designer over two weeks.
@@
Controllability: Scene can generate 3D scenes satisfying personalized demands. It achieves scene editing according to the corresponding descriptions, such as adding objects, placing objects at a location, translating the season, and so on.
@@
Diversity: Scene overcomes the constraints of conventional generation algorithms and predefined rules. By integrating and coordinating multiple subtasks, it attains a remarkable degree of flexibility and diversity in scene generation on a vast scale.
Related Works
@@
Learning Based 3D Generation. In recent years, 3D asset generation has witnessed rapid progress, combining the ideas of computer graphics and computer vision to realise the free creation of 3D content. Presently, predominant research in 3D asset generation primarily concentrate on creating individual objects (Poole et al. 2022; Liu et al. 2023; Lin et al. 2023a; Chen et al. 2023; Li et al. 2024b; Hu et al. 2024a), 3D avatars (Zhang et al. 2023a; Kolotouros et al. 2024; Hong et al. 2022; Jiang et al. 2023b), and 3D scenes (Höllein et al. 2023; Zhang et al. 2024b; Fridman et al. 2024; Zhang et al. 2023c, 2024a). Among these, ZeroShot123 (Liu et al. 2023) proposes a method based on a diffusion model, which realizes the 3D model of the target based on a picture. DreamFusion (Poole et al. 2022) proposes a NeRF-based approach that allows the model to generate a corresponding 3D model based on the input text. Compared with single object generation, it is more practical and challenging to generate large-scale scenes, including the generation of natural landforms (Hao et al. 2021; Liu et al. 2021; Chen, Wang, and Liu 2023) and borderless cities (Lin et al. 2023b; Li et al. 2023; Xie et al. 2024). CityDreamer (Xie et al. 2024) builds vast, large-scale 3D cities based on the layout of real cities, enhancing the accuracy and stability of urban reconstruction. SceneDreamer (Chen, Wang, and Liu 2023) proposes a method to generate 3D borderless scenes within 2D plots using BEV representations.
@@
Procedural Based 3D Generation. Researchers have delved into the procedural generation of natural scenes (Gasch et al. 2022; Zhang et al. 2019) and urban scenes (Lipp et al. 2011; Talton et al. 2011; Vanegas et al. 2012; Yang et al. 2013) using Procedural Content Generation (PCG). For instance, PMC (Parish and Müller 2001) proposes a procedural way to generate cities based on 2D ocean or city boundaries. It employs mathematical algorithms to generate blocks and streets and utilizes tailgating technology to generate the geometry of buildings. While these traditional computer graphic methods can produce high-quality 3D data, all parameters must be pre-entered into the procedurally generated process. This significantly constrains flexibility and practical usability. Infinigen (Raistrick et al. 2023; Zhang et al. 2024c) introduces a technique for procedurally generating realistic 3D natural objects and scenes. Although Infinigen generates infinitely assets, users are unable to customize the generated outcomes because of their specific requirements. In this paper, we propose a more convenient method to produce procedural assets.
@@
LLM Agents. Benefiting from knowledge hidden in the large-language model (LLMs) (Raffel et al. 2020; Achiam et al. 2023a; Chowdhery et al. 2023; Kenton and Toutanova 2019; Bubeck et al. 2023), researchers explore LLMs to address intricate tasks beyond canonical language processing domains. These tasks encompass areas such as mathematical reasoning (Imani, Du, and Shrivastava 2023; Wei et al. 2022), medicine (Yang et al. 2023; Jeblick et al. 2023), and planning (Zhang et al. 2023b; Gong et al. 2023; Huang et al. 2023, 2022). Thanks to powerful reasoning and generalization capabilities, LLMs act as practiced planners for different tasks. For example, (Huang et al. 2022) utilizes the expansive domain knowledge of LLMs on the internet and their emerging zero-shot planning capabilities to execute intricate task planning and reasoning. (Gong et al. 2023) investigates the application of LLMs in scenarios involving multi-agent coordination, covering a range of diverse task objectives. (Zeng et al. 2022) presents a modular framework that employs structured dialogue through prompts among multiple models. Moreover, specialized LLMs for particular applications have been explored, such as HuggingGPT (Shen et al. 2024) for vision perception tasks, VisualChatGPT (Wu et al. 2023) for multi-modality understanding, Voyager (Wang et al. 2023) and (Zhu et al. 2023), SheetCopilot (Li et al. 2024a) for office software, and Codex (Chen et al. 2021a) for Python code generation. Inspired by existing works, we explore the LLM agent to PCG software, e.g., Blender, to provide automatic 3D assets generation.
| Capability | Num. PCG | Capability | Num. PCG |
| Terrain | 12 | Water | 4 |
| Weather | 15 | Snow | 3 |
| Vegetation | 65 | Assets placement | 13 |
| Buildings | 37 | Materials | 361 |
| Blocks | 17 | Dynamic People | 4 |
| Cities | 12 | Dynamic Vegetation | 23 |
| People | 6 | Dynamic Vehicles | 3 |
Scene: All-in-One PCG Solution
We present the Scene framework for versatile scene generation, which includes PCGHub and PCGPlanner. PCGHub integrates a vast array of procedural modules and 3D assets, offering extensive procedural capabilities, while PCGPlanner coordinates these procedural modules within a well-defined algorithmic structure. The framework seamlessly incorporates user-provided PCG and assets into the workflow. Leveraging user text input, the framework enables the efficient and precise generation of diverse, high-quality scenes.
PCGHub
For scene generation tasks, the diversity of scenes is intrinsically linked to the diversity of assets. Therefore, we introduce PCGHub, a platform that integrates a wide range of PCG modules and 3D assets. It provides detailed documentation and APIs for swift integration of varied PCG techniques, addressing the limitations of traditional methods and improving content realism. A summary of the extensive PCG modules available in PCGHub is provided in Table 1. Further details are available in the supplement.
Procedural Assets. To develop PCGHub, we assemble a comprehensive collection of 172 procedural assets, categorized into natural environments, architecture, biology, environmental impacts, and science fiction elements. These assets are generated using Blender node graphs, which allow the creation of a virtually infinite range of variations through adjustments to geometry and shader parameters. From the original 2,362 parameters, we extract 263 human-interpretable parameters and encapsulate them into APIs. Each API is comprehensively documented, including functional descriptions, parameter specifications, and types, to facilitate their utilization via language models. Fig. 2 illustrates an example of API documentation, the corresponding APIs, and the resulting trees generated through these APIs. To further address the issue of limited PCG asset variety, we also collect 11,284 high-quality 3D static assets, enhancing the overall diversity and richness of the available assets.
Procedural Layout Generators. Complex patterns of arrangement and placement are inherent in various environments. To replicate these patterns, we predefine five types of layouts: scatter layout, which distributes assets randomly within a given area; grid layout, which arranges objects in a uniform grid pattern; linear layout, which arranges objects sequentially along a defined path, such as roads, rivers, or railways; nested layout, which organizes objects within a larger structure, such as buildings in a small neighborhood or attractions within a park; and area filling layout, which places objects according to specific rules to fill an entire designated area. For each layout type, we provide one or more corresponding procedural layout generators. By utilizing various objects for each layout type, we can generate a diverse range of scenes. Fig. 3 provides an example of the scatter layout, showcasing the effectiveness and diversity of procedural layout methods.
PCGPlanner
PCGPlanner utilizes the resources provided by PCGHub for efficient scene generation. As illustrated in Fig. 4, the fully automated scene generation process comprises the following stages:(1) Scene Decomposition: Analyzes the scene requirements to identify the necessary assets; (2) Terrain Generation: Constructs the foundational terrain and applies the appropriate materials; (3) Objects Generation & Retrieval: Involves generating or importing the assets required for the scene; (4) Asset Placement: Utilizes diverse layout types and procedural generators to arrange assets within the scene. Systematic Template. To endow the LLMs with the capability for modeling scenes, We introduce a systematic template to prompt LLMs in a scene modeling task. For each agent, the prompt has a similar structure defined as (, , , , ), where i corresponds to different subtasks . The constituents of the prompt are defined as follows:
@@
Document At each step, the agent is prompted by a knowledge document according to their task. We denote D as the collection that contains all the knowledge documentation pre-defined in PCGHub.
@@
Format F denotes the output format for each agent. To precisely and concisely convey information between agents, the output format of each agent is strictly defined.
@@
Examples To help agents respond strictly following the format, we demonstrate several examples E for reference in each prompt.
The systematic template guides LLMs to produce accurate outputs, enhancing the success rate and executability rate.
Scene Decomposition Stage. As illustrated in Fig. 4, this stage converts user textual descriptions into a rough list of objects for the target scene. The detailed object requirements and their corresponding procedural modules are outlined in the knowledge documentation . The process can be formulated as follows:
where is the user input, and represents an object with its associated modules and descriptions structured as {{module: description}, … {module: description}}.
Additionally, recognizing that user may omit essential details, we incorporate an Assets Clarity Assessment to identify and resolve ambiguities by actively querying users to obtain missing details for a complete preliminary plan.
Hyperparameter Generator. The Hyperparameter Generator (HyperparamGen) performs as an executor, which converts textual descriptions into PCG parameters and applies these parameters to APIs. Within PCGHub, each API is defined as a Python function with multiple predefined parameters. For each API, a corresponding knowledge document is provided to the Hyperparameter Generator. This document contains essential details required to generate the appropriate hyperparameters for executing the API. The hyperparameter generation process is formulated as follows:
where represents the initial hyperparameters and denotes the optimized parameters. In cases where the description may include incomplete parameter information, the Hyperparameter Generator is instructed to automatically select feasible parameters from available options based on the other parameters. The inferred parameters are utilized to execute the corresponding APIs in the Blender Python environment, completing all tasks efficiently.
Terrain Generation Stage. This stage is responsible for terrain generation within the scene. Unlike other assets, terrain lacks distinct characteristics that can be effectively identified using CLIP (Radford et al. 2021). Thus, a specialized terrain generation PCG is employed to ensure accurate and detailed control. As shown in the second panel of Fig. 4, the PCG provides extensive control over terrain attributes, including geometry and material properties. It enables precise adjustments to terrain slope, elevation, and features such as valleys, ensuring that the generated terrain meets specific scene requirements with both flexibility and accuracy.
Objects Generation & Retrieval Stage. Based on the detailed object descriptions provided by the Scene Decomposition Stage, we need to search within PCGHub for assets to directly import or suitable APIs to generate procedural assets. To achieve accurate retrieval, the Objects Generation & Retrieval Stage manages this process by encoding these descriptions into embeddings for retrieval. A pre-trained CLIP model is employed: text-to-text retrieval is used for procedural assets APIs, and text-to-image retrieval is used for 3D static assets. As depicted in the third panel of Fig. 4, each static asset in PCGHub is represented by a 768-dimensional vector derived from its rendering image. These vectors are compared with the input description embedding, and one of the top five most similar results is randomly selected based on cosine similarity and imported into the Blender scene.
@@
where represents the retrieved static asset from the asset collection . For text-to-text retrieval, the most relevant API is also selected by its functional description embedding. If the retrieval result is an asset, it will be directly imported into the scene project. On the other hand, the retrieved API is to be passed to the Hyperparameter Generator.
Assets Placement. In large-scale scene generation tasks, asset placement is inherently complex. It involves not only identifying spatial relationships between objects but also determining how these relationships should be represented. To address this, we propose a method based on asset relationships to guide asset placement. As illustrated in the fourth panel of Fig. 4, the LLM, guided by the predefined layout types, selects the appropriate layout for each object. Subsequently, the LLM defines the spatial relationships among assets in the scene. Each asset relationship pair is processed by the Hyperparameter Generator, which generates the necessary parameters for asset placement. These parameters are then utilized by the Procedural Layout Generators to arrange the assets accordingly in Blender.
Experiments
@@
The goals of our experiments are threefold: (i) to verify the capability of Scene for generating photorealistic large-scale scenes, including nature scenes and cities, (ii) to demonstrate the effectiveness of Scene for personalized editing, such as adding or changing,(iii) to compare different LLMs on the proposed benchmark.
Benchmark Protocol
@@
Metrics. We use Executability Rate (ER@1) and Success Rate (SR@1) to evaluate LLMs on our Scene. The former measures the proportion of proposed actions that can be executed, and the latter is used to evaluate action correctness (Chen et al. 2021b). Moreover, to quantify aesthetic quality, we adopt a unified judgment standard as a reference. We divide the aesthetics of generated scenes into five standards: Poor (1-2 points)/Below Average (3-4 points)/Average (5-6 points)/Good (7-8 points)/Excellent (9-10 points). We enlisted 35 volunteers to assess the quality of our generation, including 5 PCG experts. We compute average score (AS) and average expert score (AES) to evaluate the effectiveness of our method.
Main Results